Accelerator Update for June 16, 2008 - June 18, 2008
Monday June 16
The day shift began with Operations monitoring Tevatron (TeV) store 6226 (128.35E30) and the Recycler (RR) stash (222.2E10), with the Antiproton Source stacking (22.5E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) and the Meson MTest experiments taking beam.
At 10:10 AM, a Booster RF station (BRF7) tripped off. It reset.
At 10:36 AM, BRF3 tripped off. It reset.
At 10:46 AM, a Linac Klystron RF station (KRF5) tripped off for the first of three times in about 45 minutes. It reset.
At 11:01 AM, Operations halted stacking for about 20 minutes to let Booster experts switch out a power supply (GMPS#1). The experts realigned the system to work on three power supplies. Operations put Pbar into standby during this work.
A MI expert began a one-hour coalescing study at 1:50 PM.
Operations terminated store 6226 at 2:35 PM.
Water experts reported around 4:45 PM that they were beginning to make up water for the MI and TeV LCW systems. They also started a several day process of turning on the backup LCW pumps around the TeV in preparation for turning off the main pumps so they could clean heat exchangers (HE).
A controls expert reported at 2:49 PM that he had repaired TeV CAMAC crate E1.
Operations put the TeV into shot setup at 2:52 PM.
At 3:20 PM, BRF7 tripped off. It reset.
At 4:22 PM, LRF1 tripped off for the first of two times in two minutes. It reset.
Operations reported at 4:26 PM that the Recycler had stashed antiprotons four times during the day shift.
Operations established store 6227 at 4:44 PM with an initial luminosity of 287.0E30.
A Recycler expert began a 20-minute damper measurement at MI-60 around 6:12 PM.
At 6:38 PM, LRF1 tripped off. It reset.
At 10 PM, Operations reported that a Linac bulk power supply (L:Q7BULK) tripped off for the first of two times in two minutes. It reset.
At 11:38 PM, Operations reported that MI was aborting the beam more than usual due to bad beam positions.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Tuesday June 17
The midnight shift began with Operations monitoring store 6227 (111.95E30) and the stash (210.6E10), with Pbar stacking (18.4E10), and with BNB taking beam.
At 2:35 AM, Operations reported that the Recycler was having VSA problems.
At 6:03 AM, LRF5 tripped off. It reset.
Operations established beam to MTest at 6:16 AM.
Operations terminated store 6227 at 7:33 AM.
LCW experts reported at 7:36 AM that they were reconfiguring the HE at TeV sector E3.
Operations reported at 7:58 AM that the Recycler had stashed five times during the midnight shift.
An instrumentation expert reported at 8:38 AM that he was rebooting the MI and Recycler DCCTs.
Operations established store 6228 at 9:42 AM with an initial luminosity of 288.45E30.
At 11:08 AM, BRF3 & 8 tripped off. They reset.
At 11:22 AM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to also trip off. Everything reset.
At 3:25 PM, a Booster kicker (B:MKS10) tripped off. A person called and admitted that he had brushed against the flow switch. Operations reported that it would take about 15 minutes for the kicker to time in and turn on. At the same time, LRF2 tripped off. It reset.
At 3:27 PM, LRF2 tripped off again. It reset.
Booster resumed sending beam at 3:38 PM.
Operations reported at 4:06 PM that the Recycler had stashed four times during the day shift.
Operations reported at 5:34 PM that experts had repaired a coolant leak in the insulating gas system for the Pelletron.
Operations disabled the MTest beam at 6:17 PM. This completed the T979 experiment’s data taking.
At 8:24 PM, MIRF18 tripped off. It reset.
At 9:38 PM, the MI interlocks dropped and the stash was lost. Operators investigated and reported the following: The MI-3 enclosure interlocks dropped with no initial indication as to why, then the electrical safety system (ESS) dropped, which caused the stash to abort and the MI power supply VCBs opened.
Operations reported at 11:47 PM that the search and secure box number 11 in MI-3 would not make up. An interlock expert came in and replaced a relay. Operators searched and secured MI-3.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Wednesday June 18
The midnight shift began with Operations monitoring store 6228 (70.15E30), with no stash, with Pbar stacking (23.1E10), and with BNB off due to MI problems.
At 12:24 AM, Operations reported that they could not resume sending beam to BNB due to a problem with a power supply (E:HV860).
At 12:34 AM, a P1 transfer line dipole magnet power supply (HV703) tripped off. It reset.
Pbar resumed stacking at 12:50 AM.
BNB resumed taking beam at 1:04 AM.
At 1:10 AM, a power supply (HV860) tripped off on a current error. It reset.
At 3:27 AM, the Klystron Vernier tripped off. It reset.
Operations reported at 5:46 AM that the Pelletron had been on for 40 minutes and had not cooled the beam. Operations called in an expert.
At 5:53 AM, Operations reported that the Recycler VSA had a huge energy error.
A Recycler expert had tuned the machine, fixed the energy error, and had the Pelletron cooling antiprotons by 6:14 AM.
At 7:49 AM, KRF5 tripped off. It reset.
Operations reported at 7:58 AM that the Recycler had stashed 5 times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store. Operations will terminate store 6228 around 11 AM. D0 and CDF will make a 1.5-hour access. The next store will go in around 6 PM.
Comments and Suggestions
|last modified 06/18/2008 email Fermilab|
|Security, Privacy, Legal|